<?xml version="1.0"?>
<rss xmlns:media="http://search.yahoo.com/mrss/" version="2.0">
							<channel><title>New Posts For Thread: Dice rolling probability question</title><link>https://community.betfair.com/general_betting/go/thread/view/94082/23481970/dice-rolling-probability-question</link><description>A die is rolled 6 times. What are the probabilities that there are:a) mostly odd numbersb) mostly even numbersc) 3 of each</description><item><title>Mr Ed lets not split hairs whats your nap today ......</title><link>https://community.betfair.com/general_betting/go/thread/view/94082/23481970/dice-rolling-probability-question?post_id=419433694#419433694</link><description>Mr Ed lets not split hairs whats your nap today ......</description><pubDate>Tue, 22 Dec 2009 05:25:15 -0600</pubDate></item><item><title>a) 34.375 %b) 34.375 %c) 31.25 %total 100%</title><link>https://community.betfair.com/general_betting/go/thread/view/94082/23481970/dice-rolling-probability-question?post_id=419433690#419433690</link><description>a) 34.375 %b) 34.375 %c) 31.25 %total 100%</description><pubDate>Tue, 22 Dec 2009 05:21:40 -0600</pubDate></item><item><title>You cannot have more than 100% of anything ..... 100% means you got the lot .</title><link>https://community.betfair.com/general_betting/go/thread/view/94082/23481970/dice-rolling-probability-question?post_id=419433686#419433686</link><description>You cannot have more than 100% of anything ..... 100% means you got the lot .</description><pubDate>Tue, 22 Dec 2009 04:57:01 -0600</pubDate></item><item><title>Mande your odds add up to ...101%</title><link>https://community.betfair.com/general_betting/go/thread/view/94082/23481970/dice-rolling-probability-question?post_id=419433682#419433682</link><description>Mande your odds add up to ...101%</description><pubDate>Tue, 22 Dec 2009 04:54:10 -0600</pubDate></item><item><title>34.4, .. 34.4 .. 31. 2.. is correct .. why it adds upto 100%</title><link>https://community.betfair.com/general_betting/go/thread/view/94082/23481970/dice-rolling-probability-question?post_id=419433678#419433678</link><description>34.4, .. 34.4 .. 31. 2.. is correct .. why it adds upto 100%</description><pubDate>Tue, 22 Dec 2009 04:50:05 -0600</pubDate></item><item><title>Zola's Back Heel 21 Dec 22:04 with maths. Made me laugh.</title><link>https://community.betfair.com/general_betting/go/thread/view/94082/23481970/dice-rolling-probability-question?post_id=419433674#419433674</link><description>Zola's Back Heel 21 Dec 22:04 with maths. Made me laugh.</description><pubDate>Tue, 22 Dec 2009 04:26:32 -0600</pubDate></item><item><title>Let us know what mark you get.</title><link>https://community.betfair.com/general_betting/go/thread/view/94082/23481970/dice-rolling-probability-question?post_id=419433618#419433618</link><description>Let us know what mark you get.</description><pubDate>Mon, 21 Dec 2009 15:37:37 -0600</pubDate></item><item><title>was struggling to get my head round it. many thanks</title><link>https://community.betfair.com/general_betting/go/thread/view/94082/23481970/dice-rolling-probability-question?post_id=419433602#419433602</link><description>was struggling to get my head round it. many thanks</description><pubDate>Mon, 21 Dec 2009 15:12:56 -0600</pubDate></item><item><title>the probability of any combination of odd and even numbers over 6 throws is 1/64 or (1/2)^6, i.e. 6 consecutive throws of a dice with the outcome being a 50/50 chance (odd/even)you need to then know or work out the binomial coefficients for the diffe</title><link>https://community.betfair.com/general_betting/go/thread/view/94082/23481970/dice-rolling-probability-question?post_id=419433598#419433598</link><description>the probability of any combination of odd and even numbers over 6 throws is 1/64 or (1/2)^6, i.e. 6 consecutive throws of a dice with the outcome being a 50/50 chance (odd/even)you need to then know or work out the binomial coefficients for the diffe</description><pubDate>Mon, 21 Dec 2009 15:09:26 -0600</pubDate></item><item><title>a) = (6C4 + 6C5 + 6C6)/(2^6)= (15 + 6 + 1)/64= 22/64 = 11/32b) P(odd) = P(even) therefore the answer is the same (11/32)c) P(3 of each) = 1 - a) - b)= 1 - 22/32= 10/32= 5/16</title><link>https://community.betfair.com/general_betting/go/thread/view/94082/23481970/dice-rolling-probability-question?post_id=419433594#419433594</link><description>a) = (6C4 + 6C5 + 6C6)/(2^6)= (15 + 6 + 1)/64= 22/64 = 11/32b) P(odd) = P(even) therefore the answer is the same (11/32)c) P(3 of each) = 1 - a) - b)= 1 - 22/32= 10/32= 5/16</description><pubDate>Mon, 21 Dec 2009 15:07:47 -0600</pubDate></item></channel></rss>
